# ShortcutsBench: A Large-Scale Real-world Benchmark for API-based Agents

### v2

@@ 3.2 Dataset Construction
(“a” in Figure 2) include the “API description” (“a.4”), “API name” (“a.1”), “parameter names” (“a.2”), “parameter types” (“a.1”), “default value” (“a.2”), “return value type” (“a.3”), and “return value name” (“a.3”). The field names in [...] in Figure 2 more details about \${filename}.actionsdata, \${filename}.intentdefinition, and WFActions.json, please refer to the Appendix A.2. In existing benchmarks / datasets (Table 1), the “parameter types” (“a.1”) and “return value types” (“a.3”) are composed of primitive data types such as int and string. In addition to primitive data types, APIs in ShortcutsBench also include “enum” or “advanced data types”. Enum is composed of “the class name” and “the possible value”, with each value equipping a “value name”. We also provide the agent with a description of the “enum” in the API information. Advanced data types, such as the model (“a.1”) in app chatGPT (OpenAI, 2024), include three String types named identifier, title, and subtitle. We can comprehend them through their “type name” and “type description”.
@@
A query is a user command, such as “Tell me what the weather will be like tomorrow.” The action sequence (aka. shortcut) is the series of API calls to complete the query, with each API call referred to as an action. The action sequence identifies the steps needed to complete a query. As shown in Figure 1.b, Existing benchmarks / datasets (Table 1) collect APIs first and then use them, either fully automatically or semi-automatically, to construct query and action sequences through LLMs. In contrast, action sequences in ShortcutsBench are all human-annotated (①②③ in Figure 1). The shortcut developers are our annotators. APIs in ShortcutsBench (④ in Figure 1) are also all real-world.
Generating queries. As shown in Figure 1, existing works construct query and action sequences based on available APIs. In contrast, we construct queries based on existing action sequences and APIs. When constructing a query for a specific action sequence, we need to understand the functional description of the action sequence (“e” in Figure 2) and detailed information about the involved APIs (“a” in Figure 2). With this information, we can generate higher-quality queries. To ensure the quality of the generated queries, we also leverage the unique advantage of shortcuts: the natural language workflow descriptions (“b.7” / “c” in Figure 2). By inputting these intuitive natural language descriptions into an LLM, we can generate more accurate queries. When generating queries, we also require the model to naturally include primitive data type parameters and enum data type needed for API calls in generated queries. This helps us evaluate the agent’s ability to fill in primitive parameters in Section 3.3.2. To ensure the quality of generated queries, we use the state-of-the-art LLM, GPT-4o (OpenAI, 2024b), to generate the queries. The prompt templates we used to generate queries can be found in the Appendix A.2.
@@ 3.3 Task Definition and Metrics
We aim to address $3$ research questions regarding the performance of existing agents built using leading LLMs on ShortcutsBench with varying difficulties: (1) How do they perform in API selection? (2) How do they handle API parameter value filling, including parameters for primitive data types, enums, and outputs from previous actions? (3) Can they recognize when input is required for tasks that need system or user information?
@@
Prepare available APIs for each query. For each query $q_{i}$, we provide the LLM with a certain number of usable APIs to simulate real-world scenarios where APIs can be input into the LLM’s context. Following existing work (Meta, 2024; Qin et al., 2023; Tang et al., 2023; Xu et al., 2023; Schick et al., 2024; Hao et al., 2024), we equip each $q_{i}$ with a specific number of APIs. For each $aseq_{i}$, let $\left|APIs_{i}\right|$ represent the number of APIs involved. In addition to these $\left|APIs_{i}\right|$ APIs, we equip each query with extra APIs calculated as $max(min(x\times\left|APIs_{i}\right|,20-\left|APIs_{i}\right|),0)$, where $x\in\{3,4,5\}$. We do this because it is impractical to input all APIs into the context simultaneously. When dealing with a large number of APIs, additional retrieval is often required (Qin et al., 2023), which we do not consider in this work.
Further Processing. Considering the context limitations of LLMs, we excluded shortcuts longer than $30$ and parts using the API is.workflow.actions.runworkflow to call other shortcuts. While these shortcuts remain in our open-source dataset, they will not be included in the subsequent evaluation. We aim to study the performance of agents on queries of varying difficulties. As shown in Table 2, we categorize ShortcutsBench into $4$ difficulty levels and $8$ task types based on $\left|aseq_{i}\right|$ and “shortcut type” (Section 3.1), respectively. For more details, please refer to the Appendix A.3. When calculating the length, for branching actions like is.workflow.actions.conditional, we consider the longest branch as the length. Additionally, we ignore the lengths of looping actions like is.workflow.actions.repeat.count and special actions such as is.workflow.actions.comment. Due to the presence of branching actions, the average number of APIs involved when $p=1$ is greater than one, specifically $1.17$. For a detailed process, please refer to the Appendix A.3. The number of shortcuts in each level is denoted as $n_{p}$. Each query and action sequence is referred to as $q_{p,i}$ and $aseq_{p,i}$, with $1\leq p\leq 4$ and $1\leq i\leq n_{p}$.
@@ 3.3.1 Performance about API Selection
Following existing work (Huang et al., 2023; Patil et al., 2023; Li et al., 2023b; Xu et al., 2023; Schick et al., 2024; Hao et al., 2024), we use the accuracy of API selection as the metric. The accuracy is calculated as the number of correct API selections $m_{p}$ divided by $n_{p}$. Specifically, each time we predict an action $b_{j},1\leq j\leq|aseq_{i}|$, we provide the agent with all the correct historical actions $\{a_{1},a_{2},...,a_{j-1}\}$. We then require the agent to predict the next action. All actions predicted by the agent form the prediction sequence $bseq_{p,i}$. This method is similar to the next token prediction (NTP) in LLMs, effectively preventing a cascade of errors in subsequent action predictions due to a single incorrect prediction. During the prediction, when encountering special actions such as branching and looping, we skip predicting these actions and directly add them to the historical actions. For more details, please refer to Appendix A.4. We chose API selection accuracy over the final result for the following two additional reasons:
  • ShortcutsBench contains numerous APIs such as opening the “All Shortcuts Folder” in the Shortcuts app that do not have a return value. This makes it challenging to evaluate using existing metrics that measure the success rate of solving queries (Qin et al., 2023; Tang et al., 2023; Xu et al., 2023; Schick et al., 2024; Xu et al., 2023; Hao et al., 2024).
@@ 3.3.2 Effectiveness of API Parameter Value Filling
  • Outputs from Previous Actions: An action may either have no output or, if it does have an output, the output may be used by the following actions. In shortcuts, In ShortcutsBench, outputs that are difficult to input directly into the LLM are represented by a unique identifier (UID) and an output name (OutputName), which can be input into the LLM for processing. The agent should have the ability to correctly use the output values of previous actions.
@@ 3.3.3 Recognition of Need for Input
In this section, we aim to investigate the ability of existing API-based agents to ask systems or users for necessary information to resolve the missing information. This missing information can come from the system like clipboard (Clipboard), input files (ExtensionInput), and the current date (CurrentDate) or from the user (Ask(Apple-Inc., 2024a). For example, a parameter named tags is usually represented in a shortcut as "tags":{"Value":{"Type": "Ask"}}, where "Type": "Ask" indicates that the parameter will prompt the user for input. For more details, please refer to Appendix A.6. We use the proportion of correctly identified parameters to evaluate the agent’s ability to recognize the need for input from the system or the user. Let $n_{s}$ be the number of queries, $aska_{i}$ be the number of times the need from the system or the user appears in $aseq_{i}$, $askt_{i}$ be the number of times the need from the system or the user appears in $bseq_{i}$. The accuracy of ask for necessary information can be calculated as $Acc_{afni}=askt_{i}/aska_{i}$.
@@ 4.1 Setup
Model. Referencing existing work (Huang et al., 2023; Qin et al., 2023; Patil et al., 2023; Tang et al., 2023; Li et al., 2023b; Xu et al., 2023; Zhuang et al., 2024; Schick et al., 2024; Hao et al., 2024), considering the performance of mainstream LLMs, we selected and tested $9$ most advanced LLMs to construct API-based agent. The chosen model including $4$ closed-sourced LLMs like Gemini-1.5-Pro (DeepMind, 2024), Gemini-1.5-Flash (DeepMind, 2024), GPT-3.5-turbo (OpenAI, 2024b), and ChatGLM-4-Air (ChatGLM, 2024), and $5$ open-source LLMs like LLaMA-3-70B (Meta, 2024), QWen-2-70B (Qwen, 2024b), QWen-2-57B Qwen (2024a), Deepseek-2-Chat (236B) (DeepSeek, 2024b), and Deepseek-2-coder (236B) (DeepSeek, 2024a). Among them, Gemini-1.5-Pro, LLaMA-3-70B, QWen-2-70B, Deepseek-2-chat, and Deepseek-2-coder are LLMs benchmarked against GPT-4 OpenAI (2024a), while Gemini-1.5-Flash, ChatGLM-4-Air, and QWen-2-57B are benchmarked against GPT-3.5-turbo performance. We did not evaluate smaller LLMs like LLaMA-3-8B (Hugging Face, 2024) or Vicuna-7b-v1.5 (LMSYS, 2024) because we found that agents built on them can only handle simple tasks such as single API selection (Huang et al., 2023) and they cannot handle well on advanced tasks like parameter filling. Agents built with such models often fail to produce the required JSON actions correctly and frequently generate nonsensical outputs.
Prompt Template. Following existing work (Huang et al., 2023; Qin et al., 2023; Patil et al., 2023; Tang et al., 2023; Li et al., 2023b; Xu et al., 2023; Zhuang et al., 2024; Schick et al., 2024; Hao et al., 2024), we slightly modified the ReACT (Yao et al., 2022) templates to construct the API-based agents. Specifically, we added prompts related to shortcuts, such as the types of fillable parameters and the meanings of special statements like branches. For all $3$ research questions (RQs), we use the same prompt templates. An agent should correctly select APIs, fill in parameters, and be aware of the need to request necessary information from the system or user at appropriate times. Please refer to Appendix A.7 for more details.
@@ 4.2 Result Analysis
Figure 5: The API selection accuracy of each task type on $9$ API-based agents.
@@
  • Agents built using open-source LLMs now perform comparably to closed-source models on lower-difficulty tasks but still lag on higher-difficulty tasks. From Figure 4 we know that open-source LLMs $>=70$B match the performance of closed-source LLMs from the first $3$ difficulty tasks, significantly outperforming GPT-3.5-turbo. However, they still lag behind closed-source LLMs in handling complex tasks at the $4$-th level. Moreover, the price of open-source LLMs is significantly lower than that of GPT-3.5-turbo. For more details, please refer to Appendix A.8.
  • Existing LLM-based agents still perform poorly on tasks requiring multi-step reasoning, even Gemini-1.5-Pro level LLMs struggle with high-difficulty tasks. From Figure 4 we know that almost all LLMs handle well in API selection tasks at the level of (0,1], but only more advanced models like Gemini-1.5-Pro and QWen-2-72B can do well in higher-difficulty tasks of (1,5]. As tasks become more complex, the accuracy drops sharply. The average accuracy dropped by $19\%$ as task difficulty rose from (0,1] to (1,5], ranging from a $9\%$ decrease (Deepseek-2-chat) to a $44\%$ (ChatGLM-4-Air). From (0,1] to (5,15], accuracy fell by $46\%$, with drops from $38\%$ (Gemini-1.5-Pro) to $58\%$ (ChatGLM-4-Air).
  • Agents built with the same LLM show significant performance variations across different types of tasks. From Figure 5 we know that the performance difference of agents built with different LLM ranges from $18.04$% (ChatGLM-4-Air) to $36.70$% (Gemini-1.5-Pro).
@@
  • Compared to the API selection, for existing most intelligent LLM like Gemini-1.5-Pro, increased task difficulty has a much smaller impact on the accuracy of parameter filling, especially on using outputs from previous actions. As shown in Figure 6(a), the precision of API parameter filling of the existing most intelligent LLM like Gemini-1.5-Pro and QWen-2-72B remains similar across tasks of varying difficulty in both the upper and lower figures. This indicates that the greatest limitation of existing API-based agents in addressing user queries lies in the reasoning and planning capabilities implied by API selection.
  • Compared to API selection, the performance of API parameter filling remains a bottleneck for existing cost-effective LLMs like GPT-3.5-turbo and ChatGLM-4-Air. As shown in Figure 6(a), the performance of these LLMs in API parameter filling significantly decreases as task difficulty increases.
  • Compared to using the outputs of previous actions, extracting relevant parameters from the user’s query and filling them according to the query and API description is more challenging. As shown in Figure 6(a), the colors in the top plot (filling primitive data types and enum data types) are generally lighter than those in the bottom plot (filling the outputs of previous actions as parameters). The accuracy drop ranges from $2.55$% (GPT-3.5-turbo) to $15.39$% (Deepseek-2-Chat).
  • For existing cost-effective LLMs like GPT-3.5-turbo and ChatGLM-4-Air, errors mainly stem from incorrect output formats and wrong API selections. Figure 6(b) shows error types for tasks requiring outputs from previous actions. It can be seen that powerful LLMs like  Gemini-1.5-Pro rarely make format errors, whereas the most cost-effective models frequently make mistakes in both output format and API selection.
@@
Levels Gemini QWen Deeps. Deeps. LLaMA Gemini QWen GPT ChatGLM
1.5-Pro 2-72B 2-chat 2-coder 3-70B 1.5-Flash 2-57B 3.5-turbo 4-Air
(0, 1] 33.33 37.78 64.29 62.71 47.62 62.79 22.22 28.89 47.62
(1, 5] 45.95 50.40 55.50 60.08 44.08 53.99 37.24 37.70 48.06
(5, 15] 51.85 36.42 40.76 49.44 35.71 40.65 28.37 20.33 48.42
(15, 30] 46.67 25.00 27.59 43.14 22.22 44.64 8.11 17.14 48.89
Overall 46.59 41.97 47.90 55.18 49.89 40.71 30.74 30.55 48.28

### v3

@@ 3.2 Dataset Construction
(Figure 2.a) include the “API description” (a.4), “API name” (a.1), “parameter names” (a.2), “parameter types” (a.1), “default value” (a.2), “return value type” (a.3), and “return value name” (a.3). The field names in square brackets [] represent the original field name in the shortcut source file. For more details about \${filename}.actionsdata, \${filename}.intentdefinition, and WFActions.json, please refer to the Appendix A.2. In existing benchmarks / datasets, the “parameter types” and “return value types” are composed of primitive data types such as int and string. In addition to primitive data types, APIs in ShortcutsBench also include “enum” or “advanced data types”. Enum is composed of “the class name” and “the possible value”, with each value equipping a “value name”. We also provide the agent with a description of the “enum” in the API information. Advanced data types, such as the model (a.1), include three String types named identifier, title, and subtitle. We can comprehend them through their “type name” and “type description”.
@@
A query is a user command, such as “Tell me what the weather will be like tomorrow.” The action sequence (aka. shortcut) is the series of API calls to complete the query, with each API call referred to as an action. The action sequence identifies the steps needed to complete a query. As shown in Figure 1.b, existing benchmarks / datasets collect APIs first and then use them, either fully automatically or semi-automatically, to construct query and action sequences through LLMs. In contrast, action sequences in ShortcutsBench are all human-annotated. The shortcut developers are our annotators. APIs in ShortcutsBench are also all real-world. Moreover, we ensured the quality of action sequences by filtering shortcuts based on criteria such as whether the associated apps were paid, outdated, or relied on deprecated APIs. We also imported all shortcuts into the macOS Shortcuts app to verify their functionality.
Generating queries. As shown in Figure 1, existing works construct query and action sequences based on available APIs. In contrast, we construct queries based on existing action sequences and APIs. When constructing a query for a specific action sequence, we need to understand the functional description of the action sequence (Figure 2.e) and detailed information about the involved APIs (Figure 2.a). With this information, we can generate higher-quality queries. To ensure the quality of the generated queries, we also leverage the unique advantage of shortcuts: the natural language workflow descriptions (Figure 2.b.7 / Figure 2.c). By inputting these intuitive natural language descriptions into an LLM, we can generate more accurate queries. When generating queries, we also require the model to naturally include primitive data type parameters and enum data type needed for API calls in generated queries. This helps us evaluate the agent’s ability to fill in primitive parameters in Section 3.3.2. To ensure the quality of generated queries, we use the state-of-the-art LLM, GPT-4o (OpenAI, 2024), to generate the queries. The prompt templates we used to generate queries can be found in the Appendix A.2.
To ensure the quality of queries generated, follow existing work (Qin et al., 2024), we conducted a preliminary experiment using $3$ LLMs: GPT-4o, GPT-3.5, and Gemini-1.5-Pro, on a dataset of $100$ samples. The results showed that human evaluators rated GPT-4o generated queries the highest, outperforming the other $2$ LLMs. GPT-4o demonstrated superior performance by accurately capturing required parameters and providing clear query descriptions, meeting our criteria in $94/100$. This superior performance can largely be attributed to the natural language workflow descriptions.
@@ 3.3 Task Definition and Metrics
We aim to address $3$ research questions regarding the performance of existing agents built using leading LLMs on ShortcutsBench with varying difficulties: (1) How do they perform in API selection? (2) How do they handle API parameter value filling, including parameters for primitive data types, enums, and outputs from previous actions? (3) Can they recognize when input is required for tasks that need system or user input?
@@
Prepare available APIs for each query. For each query $q_{i}$, we provide the LLM with a certain number of usable APIs to simulate real-world scenarios where APIs can be input into the LLM’s context. Following existing work (Qin et al., 2024; Tang et al., 2023; Xu et al., 2024), we equip each $q_{i}$ with a specific number of APIs. For each $aseq_{i}$, let $\left|APIs_{i}\right|$ represent the number of APIs involved. In addition to these $\left|APIs_{i}\right|$ APIs, we equip each query with extra APIs calculated as $max(min(x\times\left|APIs_{i}\right|,20-\left|APIs_{i}\right|),0)$, where $x\in\{3,4,5\}$. We do this because it is impractical to input all APIs into the context simultaneously. When dealing with a large number of APIs, additional retrieval is often required (Qin et al., 2024; Qu et al., 2024), which we do not consider in this work.
Further Processing. Considering the context limitations of LLMs, we excluded shortcuts longer than $30$ and parts using the API is.workflow.actions.runworkflow to call other shortcuts. While these shortcuts remain in our open-source dataset, they will not be included in the subsequent evaluation. We aim to study the performance of agents on queries of varying difficulties. As shown in Table 3, we categorize ShortcutsBench into $4$ difficulty levels and $8$ task types based on $\left|aseq_{i}\right|$ and “shortcut type”, respectively. For more details, please refer to the Appendix A.3. When calculating the length, for branching actions like is.workflow.actions.conditional, we consider the longest branch as the length. Additionally, we ignore the lengths of looping actions like is.workflow.actions.repeat.count and special actions such as is.workflow.actions.comment. Due to the presence of branching actions, the average number of APIs involved when $p=1$ is greater than one, specifically $1.17$. For a detailed process, please refer to the Appendix A.3. The number of shortcuts in each level is denoted as $n_{p}$. Each query and action sequence is referred to as $q_{p,i}$ and $aseq_{p,i}$, with $1\leq p\leq 4$ and $1\leq i\leq n_{p}$.
@@ 3.3.1 Performance about API Selection
Following existing work (Huang et al., 2024b; Patil et al., 2024; Xu et al., 2024), we use the accuracy of API selection as the metric. The accuracy is calculated as the number of correct API selections $m_{p}$ divided by $n_{p}$. Specifically, each time we predict an action $b_{j},1\leq j\leq|aseq_{i}|$, we provide the agent with all the correct historical actions $\{a_{1},a_{2},...,a_{j-1}\}$. We then require the agent to predict the next action. All actions predicted by the agent form the prediction sequence $bseq_{p,i}$. This method is similar to the next token prediction (NTP) in LLMs, effectively preventing a cascade of errors in subsequent action predictions due to a single incorrect prediction. During the prediction, when encountering special actions such as branching and looping, we skip predicting these actions and directly add them to the historical actions. For more details, please refer to Appendix A.4. We chose API selection accuracy over the final result for the following two additional reasons:
  • ShortcutsBench contains numerous APIs such as opening the “All Shortcuts Folder” in the Shortcuts app that do not have a return value. This makes it challenging to evaluate using existing metrics that measure the success rate of solving queries (Qin et al., 2024; Xu et al., 2024; 2024).
@@ 3.3.2 Effectiveness of API Parameter Value Filling
  • Outputs from Previous Actions: An action may either have no output or, if it does have an output, the output may be used by the following actions. In ShortcutsBench, outputs that are difficult to input directly into the LLM are represented by a unique identifier (UID) and an output name (OutputName), which can be input into the LLM for processing. The agent should have the ability to correctly use the output values of previous actions.
@@ 3.3.3 Recognition of Need for Input
In this section, we aim to investigate the ability of existing API-based agents to ask systems or users for necessary input to resolve the missing information. This missing information can come from the system like clipboard (Clipboard), input files (ExtensionInput), and the current date (CurrentDate) or from the user (Ask(Apple-Inc., 2024a). For example, a parameter named tags is usually represented in a shortcut as "tags":{"Value":{"Type": "Ask"}}, where "Type": "Ask" indicates that the parameter will prompt the user for input. For more details, please refer to Appendix A.6. We use the proportion of correctly identified parameters to evaluate the agent’s ability to recognize the need for input from the system or the user. Let $n_{s}$ be the number of queries, $aska_{i}$, $askt_{i}$ be the number of times the need from the system or the user appears in $aseq_{i}$, $bseq_{i}$, respectively, The accuracy of ask for necessary information can be calculated as $Acc_{afni}=askt_{i}/aska_{i}$.
@@ 4.1 Setup
Model. Referencing existing work (Huang et al., 2024b; Qin et al., 2024; Li et al., 2023), considering the performance of existing LLMs, we selected $10$ most advanced LLMs to construct API-based agent. The chosen model includes $5$ closed-sourced and $5$ open-source LLMs, covering varying intelligence levels. Among them, Gemini-1.5-Pro, LLaMA-3-70B, QWen-2-70B, and Deepseek-2-chat/coder are LLMs benchmarked against GPT-4o-2024-05, while Gemini-1.5-Flash, ChatGLM-4-Air, and QWen-2-57B are benchmarked against GPT-4o-mini-2024-07 and GPT-3.5-turbo. We did not evaluate GPT-o1-preview/mini, GPT-4o, LLaMA-3.1-70b/405B, and QWen-2.5-72b, mainly due to limited access, high costs, and the fact that LLaMA and QWen are minor version improvements with recent releases.
We also compare our work with specialized API-calling fine-tuned LLMs, such as AgentLM (Zeng et al., 2024) and xLAM (Zhang et al., 2024), uncovering intriguing findings. Due to space limitations, the results are presented in Appendix A.9 for interested readers.
Prompt Template. Following existing work (Huang et al., 2024b; Qin et al., 2024; Tang et al., 2023; Zhuang et al., 2024), we slightly modified the ReACT (Yao et al., 2023) templates to construct the API-based agents. For all $3$ research questions (RQs), we use the same prompt templates. An agent should correctly select APIs, fill in parameters, and be aware of the need to request necessary input from the system or user at appropriate times. Please refer to Appendix A.7 for more details.
@@ 4.2 Result Analysis
Figure 5: The API selection accuracy of each task type on $10$ API-based agents.
From Figure 4, we can see that for tasks with a lower difficulty level, both less intelligent LLMs and more intelligent LLMs perform well. This is similar to the conclusion drawn from Table 1. However, for tasks with a higher difficulty level, only the more intelligent LLMs like Gemini-1.5-Pro, Deepseek, and QWen2 perform adequately.
The superiority of ShortcutsBench. Combined with Table 1, ShortcutsBench can effectively distinguishes between different levels of intelligence, making it a superior benchmark.
@@
  • Agents built using open-source LLMs now perform comparably to closed-source LLMs on lower-difficulty tasks but still lag on higher-difficulty tasks. From Figure 4 we know that open-source LLMs $>=70$B match the performance of closed-source LLMs from the first $3$ difficulty tasks, significantly outperforming GPT-4o-mini and GPT-3.5-turbo. However, they still lag behind closed-source LLMs in handling complex tasks at the $4$-th level. For more details, please refer to Appendix A.8.
  • Existing LLM-based agents still perform poorly on tasks requiring multi-step reasoning, even more intelligent LLMs like Gemini-1.5-Pro struggle with high-difficulty tasks. From Figure 4 we know that almost all LLMs handle well in API selection tasks at the level of (0,1], but only more advanced models like Gemini-1.5-Pro and QWen-2-72B can do well in higher-difficulty tasks of (1,5]. As tasks become more complex, the accuracy drops sharply. The average accuracy dropped by $19\%$ as task difficulty rose from (0,1] to (1,5], ranging from a $9\%$ decrease (Deepseek-2-chat) to a $44\%$ (ChatGLM-4-Air). From (0,1] to (5,15], accuracy fell by $46\%$, with drops from $38\%$ (Gemini-1.5-Pro) to $58\%$ (ChatGLM-4-Air).
  • Agents built with the same LLM show significant performance variations across different types of tasks. From Figure 5 we know that the performance difference of agents built with different LLM ranges from $15.94$% (GPT-4o-mini) to $36.70$% (Gemini-1.5-Pro).
@@
  • API selection and parameter filling both impact the agent’s performance. However, API selection has a greater effect. As shown in Figure 6(a), for existing more intelligent LLM like Gemini-1.5-Pro, increased task difficulty has a much smaller impact on the accuracy of parameter filling, especially on using outputs from previous actions. This indicates that the greatest limitation of existing API-based agents in addressing user queries lies in the reasoning and planning capabilities implied by API selection.
  • The performance of API parameter filling remains a bottleneck for existing less intelligent LLMs. As shown in Figure 6(a), the performance of less intelligent LLMs like GPT-4o-mini in API parameter filling significantly decreases as task difficulty increases.
  • Compared to using the outputs of previous actions, extracting relevant parameters from the user’s query and filling them is more challenging. As shown in Figure 6(a), the colors in the top plot (filling primitive data types and enum data types) are generally lighter than those in the bottom plot (filling the outputs of previous actions as parameters). The accuracy drop ranges from $2.55$% (GPT-3.5-turbo) to $15.39$% (Deepseek-2-Chat).
  • For existing less intelligent LLMs errors mainly stem from incorrect output formats and wrong API selections. Figure 6(b) shows error types for tasks requiring outputs from previous actions. It can be seen that powerful LLMs like  Gemini-1.5-Pro rarely make format errors, whereas the less intelligent models frequently make mistakes in output format and API selection.
@@
Levels Gemini QWen Deep Deep LLaMA Gemini QWen GPT GPT Chat
1.5 2 seek2 seek2 3 1.5 2 4o 3.5 GLM4
Pro 72B chat coder 70B Flash 57B mini turbo Air
(0, 1] 33.33 37.78 64.29 62.71 47.62 62.79 22.22 37.14 28.89 47.62
(1, 5] 45.95 50.40 55.50 60.08 44.08 53.99 37.24 40.55 37.70 48.06
(5, 15] 51.85 36.42 40.76 49.44 35.71 40.65 28.37 29.71 20.33 48.42
(15, 30] 46.67 25.00 27.59 43.14 22.22 44.64 8.11 38.89 17.14 48.89
Overall 46.59 41.97 47.90 55.18 49.89 40.71 30.74 36.71 30.55 48.28